OpenStudy (anonymous):

Solve the equation X + 7/8 = 4 A. x= 4 7/8 B. x=3 1/8 C. x= -3 1/8 D. -4 7/8 I'm and such at Fractions plz helpme

OpenStudy (anonymous):

Subtract 7/8 to both sides of the equation.

OpenStudy (anonymous):

4/1 - 7/8?

OpenStudy (anonymous):

\[x = 4-\frac{ 7 }{ 8 }\] Then make it a common denominator. \[x = \frac{ 32 }{ 8 }-\frac{ 7 }{ 8 }\] Which equals \[x = \frac{ 25 }{ 8 }\] In mixed number, it would be \[x=3\frac{ 1 }{8 }\] So the answer would be B

OpenStudy (anonymous):

Thanks! could you help with some more I have a lot to do :p

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

How do I close the Question?

OpenStudy (anonymous):

Or do I just add on?

OpenStudy (anonymous):

Click on the close button on the top left corner.